42. Calculate the standard deviation from the range values given in Example 3.5 using the Shewhart formula, s = R/d.

media%2F12f%2F12f416e8-6ccc-463f-a30d-d8
Answer #1

First we need calculate \bar{R}. So,

\bar{R}=\frac{0.6+0.8+0.8+0.6+0.8+1+0.8+0.7}{8}=0.7625

For n=5 we have d_{2}=2.326

So standrad deviation will be

s=\frac{\bar{R}}{d_{2}}=\frac{0.7625}{2.326}\approx 0.3278